# On Sum-Free Functions

### v1

@@ 2.3. Criteria for $f_{\text{\rm inv}}$ not to be $k$th order sum-free
For $k>0$, let
$\Delta(X_{1},\dots,X_{k})=\left|\begin{matrix}X_{1}&\cdots&X_{k}\cr X_{1}^{2}&\cdots&X_{k}^{2}\cr\vdots&&\vdots\cr X_{1}^{2^{k-1}}&\cdots&X_{k}^{2^{k-1}}\end{matrix}\right|\in\mathbb{F}_{2}[X_{1},\dots,X_{k}],$
and for $0\leq i\leq k$, let
$\Delta_{i}(X_{1},\dots,X_{k})=\left|\begin{matrix}X_{1}&\cdots&X_{k}\cr\vdots&&\vdots\cr X_{1}^{2^{i-1}}&\cdots&X_{k}^{2^{i-1}}\cr X_{1}^{2^{i+1}}&\cdots&X_{k}^{2^{i+1}}\cr\vdots&&\vdots\cr X_{1}^{2^{k}}&\cdots&X_{k}^{2^{k}}\end{matrix}\right|\in\mathbb{F}_{2}[X_{1},\dots,X_{k}].$
$\Delta(X_{1},\dots,X_{k})$ is known as the Moore determinant over $\mathbb{F}_{2}$ [15]. The following criterion was the basis of an approach in [7].
Theorem 2.2 ([5, Theorem 2], [7, §3.1]).
The multiplicative inverse function $f_{\text{\rm inv}}$ one $\mathbb{F}_{2^{n}}$ is not $k$th order sum-free if and only if there exists $v_{1},\dots,v_{k}\in\mathbb{F}_{2^{n}}$ such that $\Delta_{1}(v_{1},\dots,v_{k})=0$ but $\Delta(v_{1},\dots,v_{k})\neq 0$.
@@
There is another criterion by Carlet [5] for $f_{\text{\rm inv}}$ not to be $k$th order sum-free. We will revisit this criterion in Section 3.
@@
Recall that the set $\mathcal{K}_{n}$ is defined in (1.1) and Conjecture 1.1 states that for odd $n\geq 7$, $\mathcal{K}_{n}=\{3,4,\dots,n-3\}$. $n=13$ is the first value for which Conjecture 1.1 was not confirmed previously. We now confirm this case with a computer search.
@@
By §2.2 Results (4) and (8), we only have to show that $4,5,6\in\mathcal{K}_{13}$. To this end, for $k=4,5,6$, we search for $v_{1},\dots,v_{k}\in\mathbb{F}_{2^{13}}$, linearly independent over $\mathbb{F}_{2}$, such that $\Delta_{1}(v_{1},\dots,v_{k})=0$. Obviously we may assume $v_{1}=1$ (since $f_{\text{\rm inv}}$ sums to 00 on $\langle v_{1},\dots,v_{k}\rangle$, the subspace spanned by $v_{1},\dots,v_{k}$, if and only if it sums to 00 on $\langle av_{1},\dots,av_{k}\rangle$ for any $a\in\mathbb{F}_{2^{n}}^{*}$.) We represent $\mathbb{F}_{2^{13}}$ as $\mathbb{F}_{2}[X]/(f)$, where $f=X^{13}+X^{12}+X^{11}+X^{8}+1\in\mathbb{F}_{2}[X]$ is irreducible (in fact, primitive). The following are examples of the search results:
@@
Note. To prove $6\in\mathcal{K}_{13}$, we could also use §2.2 Result (5): $3\in\mathcal{K}_{13}$, $3\cdot 3<13$Rightarrow $3+3=6\in\mathcal{K}_{13}$.
@@
It is well known ([13, Theorem 3.52]) that $L_{E}(X)$ is a $q$-polynomial, i.e., of the form $\sum_{i=0}^{d}a_{i}X^{q^{i}}$; hence $L_{E}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is an $\mathbb{F}_{q}$-linear map with kernel $E$.
@@
    (Note that (3.2) implies $V\cap E=\{0\}$.) Then $s\leq\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V\leq s+1$.
@@
    $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V=s$ for some $V$ satisfying (3.2).
Proof.
@@
Assume $l\geq 2$, $l\mid n$, and that there is an $r$-dimensional $\mathbb{F}_{2}$-subspace $F$ of $\mathbb{F}_{2^{n}}$ such that $\sum_{0\neq v\in F}1/v=0$. Let $s=\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F$. The for each $0\leq t\leq n/l-s$, there is a $(tl+r)$-dimensional $\mathbb{F}_{2}$-subspace $F_{t}$ of $\mathbb{F}_{2^{n}}$ such that $\sum_{0\neq v\in F_{t}}1/v=0$ and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t}=t+s$. In particular,
@@
Assume that $F_{t}$ is given, where $t<n/l-s$. Since $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t}=t+s<n/l$, there exists $0\neq a\in\mathbb{F}_{2^{n}}$ such that $\text{\rm Tr}_{2^{n}/2^{l}}(av)=0$ for all $v\in\mathbb{F}_{2^{l}}F_{t}$, equivalently, $\text{\rm Tr}_{2^{n}/2^{l}}(av)=0$ for all $v\in F_{t}$. Thus $aF_{t}\subset L_{E}(\mathbb{F}_{2^{n}})$. Note that $\dim_{\mathbb{F}_{2}}aF_{t}=tl+r$ and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}(aF_{t})=\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t}=t+s$. By Lemma 3.2 (ii), there is an $\mathbb{F}_{2}$-subspace $V\subset\mathbb{F}_{2^{n}}$ such that $L_{E}(V)=aF_{t}$, $\dim_{\mathbb{F}_{2}}V=tl+r$ (hence $V\cap E=\{0\}$) and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V=t+s$ (hence $(\mathbb{F}_{2^{l}}V)\cap E=\{0\}$ by (3.3)). Let $F_{t+1}=E\oplus V$. Then $\dim_{\mathbb{F}_{2}}F_{t+1}=l+tl+r=(t+1)l+r$,
@@
Note. (3.4) is equivalent to $n\geq 2l(l+1)-5$. When $l>5$, this is equivalent to $n\geq 2l(l+1)$ since $l\mid n$.
@@
$2^{\circ}$ First assume
$l+2\leq\frac{n+5}{10.8}.$
By [7, Remark 4.4], $3,4,\dots,l+2\in\mathcal{K}_{n}$. Then by Theorem 3.3, for $23\leq r\leq l+$ 2,
@@
[…]
Since $k\in\mathcal{K}_{n}$ if and only $n-k\in\mathcal{K}_{n}$, the conjecture holds in this case.
@@
$\Bigl{(}\frac{n+5}{10.8}-3\Bigr{)}\frac{n+5}{10.8}<\frac{n+1}{2},$
which implies that $n\leq 83$. The pairs $(n,l)$ satisfying (3.4) and (3.5), where $n$ is odd and $13<n\leq 83$, are
$(21,3),\ (27,3),\ (33,3),\ (39,3),\ (45,3),\ (55,5),\ (65,5).$
For these pairs, by the argument in $1^{\circ}$, we only have to show that $3,4,\dots,l+2\in\mathcal{K}_{n}$. We already know that $3\in\mathcal{K}_{n}$.
$(n,l)=(21,3)$. By [7, Table 2], Conjecture 1.1 is true for $n=21$.
$(n,l)=(27,3)$. We have $4,5\in\mathcal{K}_{9}\subset\mathcal{K}_{27}$.
$(n,l)=(33,3)$. We have $4,5\in\mathcal{K}_{11}\subset\mathcal{K}_{33}$.
$(n,l)=(39,3)$. By Theorem 3.1, $4,5\in\mathcal{K}_{13}\subset\mathcal{K}_{39}$.
$(n,l)=(45,3)$. We have $4,5\in\mathcal{K}_{9}\subset\mathcal{K}_{45}$.
$(n,l)=(55,5)$. We have $4,5,6,7\in\mathcal{K}_{11}\subset\mathcal{K}_{55}$.
$(n,l)=(65,5)$. By Theorem 3.1, $4,5,6,7\in\mathcal{K}_{13}\subset\mathcal{K}_{65}$. ∎
@@
In [5], Carlet established the following criterion for $f_{\text{\rm inv}}$ not to be $k$th order sum-free using the indicator function of an $(n-k)$-dimensional subspace of $\mathbb{F}_{2^{n}}$:
Theorem 3.8 ([5]).
The multiplicative inverse function $f_{\text{\rm inv}}$ on $\mathbb{F}_{2^{n}}$ is not $k$th order sum free if and only if there exist $u_{1},\dots,u_{k}\in\mathbb{F}_{2^{n}}$, linearly independent over $\mathbb{F}_{2}$, such that
@@
the sum at the left hand side has at least $l-1$ carries. Consequently,
@@
Combining (3.7) – (3.9) gives
@@
Assume $k\geq 2$ and $2^{b_{1}}+\cdots+2^{b_{k}}=2^{n}$. Then $\min\{b_{i}:b_{i}\neq-\infty\}\geq n-k+1$ and the equality holds if and only if the multiset $\{b_{1},\dots,b_{k}\}$ equals $\{n-k+1,n-k+1,n-k+2,\dots,n-1\}$. Thus by (3.10),
@@
where $b_{i}^{\prime}=b_{i}-(n-k+1)$. ($-\infty-(n-k+1)=-\infty$.) Note that $b_{1},\dots,b_{k}\in\{-\infty,n-k+1,\dots,n-1\}$ satisfy
@@
where $m_{(2^{c_{1}},\dots,2^{c_{l}})}(X_{1},\dots,X_{k})$ is the monomial symmetric polynomial associated to the partition $(2^{c_{1}},\dots,2^{c_{l}})$ ([14, I.2]). We shall call a partition whose parts are powers of $2$ a $2$-adic partition. Thus,
@@
The multiplicative inverse function $f_{\text{\rm inv}}$ is not $k$th order sum-free if and only if there are $u_{1},\dots,u_{k}\in\mathbb{F}_{2^{n}}$ such that $\Delta(u_{1},\dots,u_{k})\neq 0$ but $\Theta_{k}(u_{1},\dots,u_{k})=0$.
@@
(The enumeration of all terms in $\Theta_{4}$ are given in the appendix.)
@@
Thus, (3.11) is satisfied when
@@
We may assume that $n\geq 17$. By Theorem 3.6, we may assume that $(5-1)(5+2)>(n+1)/2$, i.e., $n<55$. By Corollary 3.7, we may further assume $3\nmid n$. Therefore, we only have to consider $n=25,35$.
When $n=25$, we only have to check that $3,\dots,12\in\mathcal{K}_{25}$. We already know that $3,4,5\in\mathcal{K}_{25}$ (§2.2 Results (1), (6), and Theorem 3.12). Using §2.2 Result (5) repeatedly, we see that 6=3+3, 7=3+4,, 11=3+8formulae-sequence633formulae-sequence73411386=3+3,\,7=3+4,\,$\dots$,\,11=3+86 = 3 + 3 , 7 = 3 + 4 , … , 11 = 3 + 8 all belong to $\mathcal{K}_{25}$. By Corollary 3.4, $13=2\cdot 5+3\in\mathcal{K}_{25}$, whence $12=25-13\in\mathcal{K}_{25}$.
[…]

### v2

@@ 2.4. Criteria for $f_{\text{\rm inv}}$ not to be $k$th order sum-free
Recall that $k\in\mathcal{K}_{n}$ means that $f_{\text{\rm inv}}$ is not $k$th order sum-free. The following criterion was the basis of an approach in [7]. We follow the notation of §2.3 with $q=2$.
Theorem 2.4 ([5, Proposition 1], [7, §3.1]).
We have $k\in\mathcal{K}_{n}$ if and only if there exist $v_{1},\dots,v_{k}\in\mathbb{F}_{2^{n}}$ such that $\Delta_{1}(v_{1},\dots,v_{k})=0$ but $\Delta(v_{1},\dots,v_{k})\neq 0$.
@@
There is another criterion by Carlet [5] for $k\in\mathcal{K}_{n}$. We will revisit this criterion in Section 3.
@@
Recall that the set $\mathcal{K}_{n}$ is defined in equation (1.1) and Conjecture 1.1 states that for odd $n\geq 7$, $\mathcal{K}_{n}=\{3,4,\dots,n-3\}$. $n=13$ is the first value for which Conjecture 1.1 was not confirmed previously. We now confirm this case with a computer search.
@@
By §2.2 Results (4) and (8), we only have to show that $4,5,6\in\mathcal{K}_{13}$. To this end, for $k=4,5,6$, we search for $v_{1},\dots,v_{k}\in\mathbb{F}_{2^{13}}$, linearly independent over $\mathbb{F}_{2}$, such that $\Delta_{1}(v_{1},\dots,v_{k})=0$. Obviously we may assume $v_{1}=1$ (since $f_{\text{\rm inv}}$ sums to 0 on $\langle v_{1},\dots,v_{k}\rangle$, the subspace spanned by $v_{1},\dots,v_{k}$, if and only if it sums to 0 on $\langle av_{1},\dots,av_{k}\rangle$ for any $a\in\mathbb{F}_{2^{n}}^{*}$.) We represent $\mathbb{F}_{2^{13}}$ as $\mathbb{F}_{2}[X]/(f)$, where $f=X^{13}+X^{12}+X^{11}+X^{8}+1\in\mathbb{F}_{2}[X]$ is irreducible (in fact, primitive). The following are examples of the search results:
@@
Note. To prove $6\in\mathcal{K}_{13}$, we could also use §2.2 Result (5): $\in\mathcal{K}_{13}, 33$$\cdot 3<13$ $\Rightarrow$ $3+3=6\in\mathcal{K}_{13}$.
@@
It is well known ([16, Theorem 3.52]) that $L_{E}(X)$ is a $q$-polynomial, i.e., of the form $\sum_{i=0}^{d}a_{i}X^{q^{i}}$; hence $L_{E}:\mathbb{F}_{q^{n}}\to\mathbb{F}_{q^{n}}$ is an $\mathbb{F}_{q}$-linear map with kernel $E$.
@@
    (Note that condition (3.2) implies $V\cap E=\{0\}$.) Then $s\leq\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V\leq s+1$.
@@
    We have $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V=s$ for some $V$ satisfying condition (3.2).
Note that in Lemma 3.2, $s\leq r$, which gives Corollary 3.4 later.
Proof of Lemma 3.2.
@@
Assume $l\geq 2$, $l\mid n$, and that there is an $r$-dimensional $\mathbb{F}_{2}$-subspace $F$ of $\mathbb{F}_{2^{n}}$ such that $\sum_{0\neq v\in F}1/v=0$. Let $s=\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F$. Then for each $0\leq t\leq n/l-s$, there is a $(tl+r)$-dimensional $\mathbb{F}_{2}$-subspace $F_{t}$ of $\mathbb{F}_{2^{n}}$ such that $\sum_{0\neq v\in F_{t}}1/v=0$ and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t}=t+s$. In particular,
@@
Assume that $F_{t}$ is given, where t<n/lst<n/l-s. Since $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t}=t+s<n/l$, there exists $0\neq a\in\mathbb{F}_{2^{n}}$ such that $\text{\rm Tr}_{2^{n}/2^{l}}(av)=0$ for all $v\in\mathbb{F}_{2^{l}}F_{t}$, equivalently, $\text{\rm Tr}_{2^{n}/2^{l}}(av)=0$ for all $v\in F_{t}$. Thus $aF_{t}\subset L_{E}(\mathbb{F}_{2^{n}})$. Note that $\dim_{\mathbb{F}_{2}}aF_{t}=tl+r$ and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}(aF_{t})=\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}F_{t}=t+s$. By Lemma 3.2 (ii), there is an $\mathbb{F}_{2}$-subspace $V\subset\mathbb{F}_{2^{n}}$ such that $L_{E}(V)=aF_{t}$, $\dim_{\mathbb{F}_{2}}V=tl+r$ (hence $V\cap E=\{0\}$) and $\dim_{\mathbb{F}_{2^{l}}}\mathbb{F}_{2^{l}}V=t+s$ (hence $(\mathbb{F}_{2^{l}}V)\cap E=\{0\}$ by equation (3.3)). Let $F_{t+1}=E\oplus V$. Then $\dim_{\mathbb{F}_{2}}F_{t+1}=l+tl+r=(t+1)l+r$,
@@
Note. Inequality (3.4) is equivalent to $n\geq 2l(l+1)-5$. When $l>5$, this is equivalent to $n\geq 2l(l+1)$ since $l\mid n$.
@@
$1^{\circ}$ First assume
$l+2\leq\frac{3n-9}{13}.$
By [7, Theorem 4.2], $3,4,\dots,l+2\in\mathcal{K}_{n}$. Then by Theorem 3.3, for $3\leq r\leq l+2$,
@@
$\mathcal{K}_{n}\supset\Bigl\{tl+r:3\leq r\leq l+2,\ 0\leq t\leq\frac{n}{l}-r\Bigr\}\supset\Bigl\{3,4,\dots,\frac{n-1}{2}\Bigr\}.$
Since $k\in\mathcal{K}_{n}$ if and only if $n-k\in\mathcal{K}_{n}$, the conjecture holds in this case.
@@
$\Bigl(\frac{3n-9}{13}-3\Bigr)\frac{3n-9}{13}<\frac{n+1}{2},$
which implies that $n<27$. The pair $(n,l)$ satisfying inequalities (3.4), where $n$ is odd and $13<n<27$, is $(21,3)$. By [7, Table 2], Conjecture 1.1 is true for $n=21$. ∎
@@
In [5], Carlet established the following criterion for $k\in\mathcal{K}_{n}$ using the indicator function of an (nk)(n-k)-dimensional subspace of $\mathbb{F}_{2^{n}}$:
Theorem 3.8 ([5, §5.2.4]).
We have $k\in\mathcal{K}_{n}$ if and only if there exist $u_{1},\dots,u_{k}\in\mathbb{F}_{2^{n}}$, linearly independent over $\mathbb{F}_{2}$, such that
@@
the sum at the left hand side has at least l1l-1 carries. Consequently,
@@
Combining equations (3.7), (3.8) and inequality (3.9), we have
@@
Assume $k\geq 2$ and $2^{b_{1}}+\cdots+2^{b_{k}}=2^{n}$. Then $\min\{b_{i}:b_{i}\neq-\infty\}\geq n-k+1$ and the equality holds if and only if the multiset $\{b_{1},\dots,b_{k}\}$ equals $\{n-k+1,n-k+1,n-k+2,\dots,n-1\}$. Thus by equation (3.10),
@@
where $b_{i}^{\prime}=b_{i}-(n-k+1)$. ($\infty-(n-k+1)=-\infty$.) Note that $b_{1},\dots,b_{k}\in\{-\infty,n-k+1,\dots,n-1\}$ satisfy
@@
where $m_{(2^{c_{1}},\dots,2^{c_{l}})}(X_{1},\dots,X_{k})$ is the monomial symmetric polynomial associated to the partition $(2^{c_{1}},\dots,2^{c_{l}})$ ([17, I.2]). We shall call a partition (of a non-negative integer) a $2$-adic partition if the parts are powers of $2$. Thus,
@@
With this notation, Theorem 3.8 is reformulated as follows:
@@
We have $k\in\mathcal{K}_{n}$ if and only if there are $u_{1},\dots,u_{k}\in\mathbb{F}_{2^{n}}$ such that $\Delta(u_{1},\dots,u_{k})\neq 0$ but $\Theta_{k}(u_{1},\dots,u_{k})=0$.
@@
The enumeration of all terms in $\Theta_{4}$ are given in the appendix.
The next theorem is an application of Theorem 3.11
@@
Thus, inequality (3.11) is satisfied when
@@
We may assume that $n\geq 17$. By Theorem 3.6, we may assume that (51)(5+2)>(n+1)/2(5-1)(5+2)>(n+1)/2, i.e., $n<55$. By Corollary 3.7, we may further assume $3\nmid n$. Therefore, we only have to consider $n=25,35$.
When $n=25$, we only have to check that $3,\dots,12\in\mathcal{K}_{25}$. We already know that $3,4,5\in\mathcal{K}_{25}$ (§2.2 Results (1), (6), and Theorem 3.12). Using §2.2 Result (5) repeatedly, we see that $6=3+3,\,7=3+4,\,\dots,\,11=3+8$ all belong to $\mathcal{K}_{25}$. By Corollary 3.4, $13=2\cdot 5+3\in\mathcal{K}_{25}$, whence $12=25-13\in\mathcal{K}_{25}$.
When $n=35$, we only have to check that $3,\dots,17\in\mathcal{K}_{35}$. As in the above, we have $3,\,4,\,5,\,6(=3+3),\,\dots,\,14(=3+11)\in\mathcal{K}_{35}$. Also, $15\in\mathcal{K}_{35}$ since $\gcd(15,35)>1$. By Corollary 3.4, both $18=3\cdot 5+3$ and $19=3\cdot 5+4$ belong to $\mathcal{K}_{35}$. Therefore, $16=35-19\in\mathcal{K}_{35}$ and $17=35-18\in\mathcal{K}_{35}$. ∎